# references.bib

### v2

@@ Abstract
We propose the Block Coordinate Descent Network Simplex (BCDNS) method for solving large-scale discrete Optimal Transport (OT) problems. BCDNS integrates the Network Simplex (NS) algorithm with a block coordinate descent (BCD) strategy, decomposing the full problem into smaller subproblems per iteration and reusing basis variables to ensure feasibility. We prove that BCDNS terminates in a finite number of iterations with an exact optimal solution, and we characterize its per-iteration complexity as $O(sN)$, where $s\in(0,1)$ is a user-defined parameter and $N$ is the total number of variables. Numerical experiments demonstrate that BCDNS matches the classical NS method in solution accuracy, reduces memory footprint compared to Sinkhorn algorithm, achieves speed-ups of up to tens of times over the classical NS method, and exhibits runtime comparable to high-precision Sinkhorn algorithm.
@@ 1.   Introduction
The Optimal Transport (OT) problem is to find the most efficient way to transport resources between a finite number of locations while minimizing transportation costs. It is rooted in mathematical optimization and has a rich history dating back to the 18th century. Gaspard Monge [monge1781deblais] first proposed it in 1781. Then in the mid-20th century, Leonid Kantorovich [kantorovich1942translocation] extended Monge’s framework to encompass continuous distributions, reformulating it as a linear programming model and significantly broadening its theoretical foundation and practical applications. In this thesis, we focus specifically on solving the discrete OT problem.
The OT problem has found applications in various disciplines. In machine learning, the OT problem has been used for tasks such as domain adaptation [courty2016optimal] and generative modeling  [arjovsky2017wasserstein]. In economics, the OT problem has provided insights into resource allocation, market equilibrium, and pricing mechanisms [galichon2016optimal]. In biology, the OT problem has been applied to analyze genetic expression patterns [schiebinger2019optimal].
@@
The OT problem has found applications in various disciplines. In machine learning, the OT problem has been used for tasks such as domain adaptation [courty2016optimal] and generative modeling [arjovsky2017wasserstein]. In economics, the OT problem has provided insights into resource allocation, market equilibrium, and pricing mechanisms [galichon2016optimal]. In biology, the OT problem has been applied to analyze genetic expression patterns [schiebinger2019optimal]. In geophysics and signal processing, the OT problem has been employed for seismic waveform comparison and inversion. [liao2022fast].
@@
Several methods have been developed to solve the OT problem. Each method has distinct strengths and limitations. The Hungarian method is a classical approach for solving linear assignment problems and can also be applied to OT problems [kuhn1955hungarian]. It achieves exact optimal solutions by constructing a cost matrix and iteratively reducing reducing dimensions of the matrix. However, its time complexity is $O(N^{2})$, where $N$ is the number of variables in OT problem, and hence it is not applicable for the large-scale OT problem. The Sinkhorn algorithm is a popular method for the large-scale OT problem due to its computational efficiency [cuturi2013sinkhorn]. It solves the OT problems with regularization of entropy, which is an approximation of the OT problem. Therefore it provides approximate solutions of the OT problem, and it may not be suitable for applications that require exact solutions. In recent years, many improvements based on the Sinkhorn algorithm have been proposed, such as the Greenkhorn algorithm [altschuler2017Near] and the Adaptive Primal–Dual Accelerated Gradient Descent (APDAGD) algorithm [dvurechensky2018Computational]. However, these methods still operate within the Sinkhorn framework and do not overcome the limitation of producing only approximate OT solutions. In contrast, the Network Simplex (NS) method, a specialized variant of the simplex method, is well known to yield exact solutions. Michael D. Grigoriadis [grigoriadis1986efficient] reported that the NS method is reliable among several algorithms. However, the NS method is not a polynomial time algorithm and it takes $O(N)$ time in each iteration, which leads to slow computational performance when applied to large-scale problems. We note that the above three methods needs at least $O(N)$ space complexity. To overcome the difficulty of large space complexity for the OT problem, Yue Xie et al. [xie2023randomized] proposed the Block Coordinate Descent (BCD) method for solving the OT problem. Their method constructs a small subproblem by randomly selecting variables based on predefined rules and solves it in each iteration. Since the subproblem is small, the BCD method does not require the O(N) space complexity. However, since the BCD method proposed by Yue Xie et al. adapts the random variable selection, it is not guaranteed to find an optimal solution within a finite number of iterations. Moreover, it sometimes takes a large number of iterations to get a reasonable approximate solution.
@@
The motivation behind this research lies in the growing demand for efficient methods capable of solving large-scale OT problems with highly accurate solutions. As data sizes continue to grow, we recognize the increasing importance of solving large-scale OT problems with both speed and accuracy. To address these challenges, we aim to develop a method that enables fast computation of exact solutions for large-scale OT problems.
@@
To achieve the goal of efficiently solving large-scale OT problems, we introduce the idea of the BCD method. The BCD method divides a large-scale OT problem into smaller subproblems, enabling a reduction in computational complexity. Although the BCD method is generally considered slow when exact solutions are required, we observe that the optimal solutions of the OT problem are typically sparse—due to the relatively small number of equality constraints compared to the number of variables—which significantly improves the efficiency of BCD. By employing the NS method to solve the subproblems, we guarantee that the final solution is also exact.
In this paper, we propose the Block Coordinate Descent Network Simplex (BCDNS) method, a novel algorithm that integrates the BCD and NS methods to achieve computational efficiency and solution accuracy. Our approach reduces computational complexity by decomposing large-scale OT problems into smaller subproblems, while the NS method ensures that each subproblem is solved with precision. Numerical experiments demonstrate significant performance gains of the BCDNS method over traditional approaches, addressing key limitations of existing methods.
In the BCD method, the selection of the block coordinate is crucial. Compared to existing BCD method [xie2023randomized] that randomly selects variables based on matrix properties, we propose that the BCDNS method first selects the basis variables from the previous subproblem as the initial solution for the next subproblem. This strategy ensures that the subproblem always has a feasible and an optimal solution. Moreover, by integrating the principles of the NS method, we ensure that the BCDNS method converges to the exact optimal solution within a finite number of computations. Furthermore, we categorize variables into two groups based on their likelihood of becoming optimal basis variables. This targeted selection process prioritizes variables with a higher probability of being optimal basis variables, thereby improving the efficiency of finding the optimal solution.
The outline of this paper is as follows. In Section 2, we introduce the foundational concepts of the OT problem and review several existing methods for solving it. In Section 3, we describe the core principles of the proposed BCDNS method. We first present its prototype along with the corresponding theoretical proofs. Then we develop a concrete implementation based on the prototype. Section 4 reports numerical experiments, including parameter estimation for the BCDNS method and a comparative performance analysis against existing approaches. Finally, Section 5 concludes the paper by summarizing the key contributions and discussing potential directions for future research.
@@ 2.   Discrete Optimal Transport
We consider a discrete OT problem [villani2009optimal] between $n$ supply locations and $m$ demand locations. Let $\{1,2,\dots,n\}$ denote the index set of supply points, and $\{1,2,\dots,m\}$ denote that of demand points.
@@
    where $I_{n}$ denotes the $n\times n$ identity matrix, and $1_{n}$ is an $n$-dimensional column vector of ones.
With this notation, we can rewrite the OT problem as:
@@ 3.   Existing Methods for Optimal Transport
In this section, we review several representative algorithms for solving the discrete OT problem. These include the classical NS method, the Sinkhorn algorithm based on entropy regularization which is widely used for large-scale problems, and the recently proposed BCD method.
@@ 3.1 Network Simplex Method for Optimal Transport
As shown in Appendix A, the constraint matrix $A\in\mathbb{R}^{M\times N}$ in the OT problem has rank $M-1$. This allows us to partition the set of variable indices $\mathcal{N}=\{1,\dots,N\}$ into two subsets: a basis index set $\mathcal{B}\subseteq\mathcal{N}$ with $|\mathcal{B}|=M-1$, and its complement $\mathcal{N}\setminus\mathcal{B}$. The columns of $A$ corresponding to $\mathcal{B}$, denoted $A_{B}\in\mathbb{R}^{M\times(M-1)}$, form a full-rank submatrix. The basic variables $\mathbf{x}_{B}$ can then be uniquely determined by the system $A_{B}\mathbf{x}_{B}=\mathbf{b}$, while the non-basic variables $\mathbf{x}_{N}$ are fixed at zero. A basic feasible solution further satisfies the non-negativity constraint $\mathbf{x}_{B}\geq 0$.
To determine whether a given basic solution is optimal, we examine its reduced-cost vector. Define the dual potential vector $\boldsymbol{\pi}\in\mathbb{R}^{M}$ via the dual feasibility condition $\boldsymbol{\pi}=(A_{B}^{\top})^{-1}\mathbf{c}_{B}$, where $\mathbf{c}_{B}$ is the cost subvector corresponding to the basis $B$. Then, for any index $k\in N\setminus B$, the reduced cost is computed as $r_{k}=c_{k}-\boldsymbol{\pi}^{\top}A_{k}$. Due to the structure of the OT problem, each column $A_{ij}$ has exactly one $+1$ and one $-1$, so the reduced cost simplifies to the closed-form expression
@@
Optimality of the current basic solution is certified if all reduced costs are non-negative, i.e., $r_{ij}\geq 0$ for all $i,j$. If a variable $x_{i^{*}j^{*}}$ has $r_{i^{*}j^{*}}<0$, it indicates a direction of descent in the objective, and that variable is introduced into the basis.
To perform a pivot, we add the index $(i^{*},j^{*})$ to the basis $B$, inducing a unique cycle in the associated flow graph. Then determine the step size $\Delta$ by the minimum flow along the negative edges of the cycle, ensuring feasibility. Subsequently, update the flow: $x_{i^{*}j^{*}}\leftarrow\Delta$, and for each edge $(u,v)$ along the cycle, the value of $x_{uv}$ is incremented or decremented by $\Delta$ depending on the cycle direction. Finally, the variable that reaches zero and remove it from the basis, completing the pivot step [vanderbei2001linear].
Each iteration of the NS method involves computing the dual potential vector $\boldsymbol{\pi}$ in $O(M)$ time, evaluating all reduced costs in $O(N)$, and performing pivoting operations in $O(M)$. Since each pivot strictly improves the objective function by at least $-r_{i^{*}j^{*}}>0$, and the number of possible basic feasible solutions is finite, the NS method terminates in a finite number of iterations.
@@ 3.2 Block Coordinate Descent Method for Optimal Transport
In this subsection we introduce the BCD method for the OT problem in the LP form (2).
@@
$\𝒩D(\mathbf{x};\mathcal{H})\triangleq\arg\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}.$ (3)
$\𝒩q(\mathbf{x};\mathcal{H})\triangleq\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}.$ (4)
Here, $D(\mathbf{x};\mathcal{H})$ represents the set of optimal solutions for the subproblem with $\mathcal{H}$, and $q(\mathbf{x};\mathcal{H})$ denotes the optimal objective value corresponding to $d\in D(\mathbf{x};\mathcal{H})$ [tseng2009block].
@@
The existing BCD method [xie2023randomized], randomly selects block coordinates based on certain properties of specific matrices. In [xie2023randomized], the authors propose the Accelerated Random Block Coordinate Descent algorithm and establish the following convergence theorem.
@@
The gap between the current objective function value and the optimal objective function value converges with rate $1-v$, i.e.,
@@
These approaches provide a simple mechanism for constructing subproblems by selecting block coordinates. However, when the size n of the OT problems is large, $v$ becomes extremely small, which leads to slow convergence. Furthermore, these approaches do not account for the sparsity of the solution of the OT problem. Many variables that ultimately take zero values continue to be included in the block coordinates at each iteration. Moreover, these approaches have an inherent limitation: they cannot guarantee convergence to the optimal solution within a finite number of iterations. This limitation arises from the possibility of extreme cases where the randomly selected block of coordinates rare includes the optimal basis variables.
@@ 3.3 Sinkhorn Algorithm for Optimal Transport
Instead of solving the classical OT problem defined in (1), the Sinkhorn algorithm [cuturi2013sinkhorn] introduces an entropy regularization term
@@
The regularization parameter $\varepsilon$ introduces a smoothing effect in the optimal transport problem. As $\varepsilon$ decreases, the solution becomes closer to the exact OT plan, but the algorithm may converge more slowly. Conversely, larger values of $\varepsilon$ result in faster computations, but with a higher approximation error in the transport plan.

### v3

# Block Coordinate Descent Network Simplex Methods for Optimal Transport

@@ Abstract
We propose the Block Coordinate Descent Network Simplex (BCDNS) method for solving large-scale Optimal Transport (OT) problems. BCDNS integrates the Network Simplex (NS) algorithm with a block coordinate descent (BCD) strategy, decomposing the full problem into smaller subproblems per iteration and reusing basis variables to ensure feasibility. We prove that BCDNS terminates in a finite number of iterations with an exact optimal solution, and we characterize its per-iteration complexity as $O(sN)$, where $s\in(0,1)$ is a user-defined parameter and $N$ is the total number of variables. Numerical experiments show that BCDNS attains the same exact objective value as the classical NS method while substantially reducing runtime on large instances, with speed-ups of up to tens of times over NS. In addition, comparisons with the stabilized Sinkhorn algorithm under different regularization levels illustrate that BCDNS converges to the unregularized OT optimum, and its runtime is competitive with Sinkhorn when high accuracy is required.
@@ 1.  Introduction
The Optimal Transport (OT) problem is to find the most efficient way to transport resources between a finite number of locations while minimizing transportation costs. It is rooted in mathematical optimization and has a rich history dating back to the 18th century. Gaspard Monge [13] first proposed it in 1781. Then in the mid-20th century, Leonid Kantorovich [10] extended Monge’s framework to encompass continuous distributions, reformulating it as a linear programming model and significantly broadening its theoretical foundation and practical applications. In this paper, we focus specifically on solving the discrete OT problem.
@@
The OT problem has found applications in various disciplines. In machine learning, the OT problem has been used for tasks such as domain adaptation [5] and generative modeling [3]. In economics, the OT problem has provided insights into resource allocation, market equilibrium, and pricing mechanisms [8]. In biology, the OT problem has been applied to analyze genetic expression patterns [14]. In geophysics and signal processing, the OT problem has been used for seismic waveform comparison and inversion [12].
@@
Several methods have been developed to solve the OT problem. Each method has different strengths and limitations. The Hungarian method is a classical algorithm for the linear assignment problem [11], which can be viewed as a special case of OT where the transport plan is restricted to a one-to-one matching. It produces an exact optimum in $O(n^{3})$ time for an $n\times n$ assignment instance. However, for general OT, the Hungarian method is no longer applicable without expanding the problem size, which becomes prohibitive for large-scale problems. The Sinkhorn algorithm is a popular method for the large-scale OT problem due to its computational efficiency [6]. It solves the OT problems with regularization of entropy, which is an approximation of the OT problem. Therefore it provides approximate solutions of the OT problem, and it may not be suitable for applications that require exact solutions. In recent years, many improvements based on the Sinkhorn algorithm have been proposed, such as the Greenkhorn algorithm [2] and the Adaptive Primal–Dual Accelerated Gradient Descent algorithm [7]. However, these methods still operate within the Sinkhorn framework and do not overcome the limitation of producing only approximate OT solutions. In contrast, the Network Simplex (NS) method [9], a specialized variant of the simplex method, is well known to yield exact solutions. However, the NS method takes $O(N)$ time in each iteration, which leads to slow computational performance when applied to large-scale problems. Yue Xie et al. [18] proposed a block coordinate descent (BCD) approach for solving large-scale OT problems. At each iteration, the method constructs a restricted subproblem by selecting a subset of variables according to a prescribed sampling rule and then solves the subproblem while keeping the remaining variables fixed. By operating on a much smaller working set, the BCD approach can substantially reduce the computational burden per-iteration and, in practice, also reduces working memory compared to methods that process all $N$ variables at once. However, since the BCD method proposed by Yue Xie et al. adapts the random variable selection, it is not guaranteed to find an optimal solution within a finite number of iterations. Moreover, it sometimes takes a large number of iterations to get a reasonable approximate solution.
@@
The motivation behind this research lies in the growing demand for efficient methods capable of solving large-scale OT problems with highly accurate solutions. As data sizes continue to grow, we recognize the increasing importance of solving large-scale OT problems with both speed and accuracy. To address these challenges, our aim is to develop a method that enables fast computation of exact solutions for large-scale OT problems.
@@
To achieve the goal of efficiently solving large-scale OT problems, we introduce the idea of the BCD method. The BCD method divides a large-scale OT problem into smaller subproblems, allowing a reduction in computational complexity. Although the BCD method is generally considered slow when exact solutions are required, we observe that the optimal solutions of the OT problem are typically sparse, which significantly improves the efficiency of the BCD method. By employing the NS method to solve the subproblems, we guarantee that the final solution is also exact.
In this paper, we develop the Block Coordinate Descent Network Simplex (BCDNS) method for solving large-scale OT problems with exact optimality. Compared to existing BCD method [18] that randomly selects variables based on matrix properties, we introduced a basis variable succession strategy to reuse basis information across iterations, which ensures monotone improvement and finite termination. Furthermore, we propose a grouped variable selection mechanism to reduce unnecessary reduced-cost evaluations. Theoretical analysis establishes convergence to the global optimum in finitely many iterations, and numerical experiments demonstrate that the proposed method significantly outperforms the classical NS method and is competitive with high-precision Sinkhorn algorithms in terms of runtime when high accuracy is required.
The remainder of this paper is organized as follows. Section 2 formulates the discrete OT problem and introduces the notation used throughout the paper. Section 3 reviews representative approaches for OT, including the NS method, the BCD methods, and the Sinkhorn algorithm. Section 4 presents the proposed BCDNS framework, together with concrete block selection strategies and theoretical properties. Section 5 reports the numerical experiments, including the test settings and the empirical results. Finally, Section 6 concludes the paper and discusses future directions.
@@ 2.  Discrete Optimal Transport
We consider a discrete OT problem [17] between $n$ supply locations and $m$ demand locations. Let $\{1,2,\dots,n\}$ denote the index set of supply points, and $\{1,2,\dots,m\}$ denote that of demand points.
@@
    where $I_{n}$ denotes the identity matrix $n\times n$, and $1_{n}$ is a column vector of dimensions $n$.
With this notation, we can rewrite the OT problem as follows:
@@ 3.  Existing Methods for Optimal Transport
In this section, we review several representative algorithms for solving the discrete OT problem. These include the classical NS method, the Sinkhorn algorithm based on entropy regularization, which is widely used for large-scale problems, and the recently proposed BCD method.
@@ 3.1 Network Simplex Method for Optimal Transport
It is well known that the node–arc incidence matrix of a connected graph has rank $\lvert V\rvert-1$ [1]. Since the OT network is the complete bipartite graph with $M=n+m$ nodes, we have $\operatorname{rank}(A)=M-1$. This allows us to partition the set of variable indices $\mathcal{N}=\{1,\dots,N\}$ into two subsets: a set of basis indices $\mathcal{B}\subseteq\mathcal{N}$ with $|\mathcal{B}|=M-1$, and its complement $\mathcal{N}\setminus\mathcal{B}$. The columns of $A$ corresponding to $\mathcal{B}$, denoted $A_{B}\in\mathbb{R}^{M\times(M-1)}$, form a full-rank submatrix. The basic variables $\mathbf{x}_{B}$ are determined by the equality constraints
$A_{B}\mathbf{x}_{B}=\mathbf{b},$
while the nonbasic variables satisfy $\mathbf{x}_{N}=\mathbf{0}$. A basic feasible solution (BFS) additionally requires $\mathbf{x}_{B}\geq\mathbf{0}$. To assess the optimality of a BFS, we examine the reduced costs derived from dual variables(node potentials). Let $\boldsymbol{\pi}\in\mathbb{R}^{M}$ denote a vector of node potentials. For a given basis $B$ (which corresponds to a spanning tree), we choose $\boldsymbol{\pi}$ to satisfy the dual optimality equations
$A_{B}^{\top}\boldsymbol{\pi}=\mathbf{c}_{B}.$ (3)
Since $A_{B}$ has rank M1M-1, the solution of (3) is unique up to an additive constant; we fix this degree of freedom by setting the potential of an arbitrary reference node to zero. In practice, $\boldsymbol{\pi}$ can be computed in $O(M)$ time by traversing the spanning tree.
Given $\boldsymbol{\pi}$, the reduced cost of any variable $\ell$ is
$\bar{c}_{\ell}\;=\;c_{\ell}-A_{\ell}^{\top}\boldsymbol{\pi},$ (4)
where $A_{\ell}$ denotes the $\ell$-th column of $A$. Equivalently, in vector form $\bar{\mathbf{c}}=\mathbf{c}-A^{\top}\boldsymbol{\pi}$. Due to the structure of the OT problem, each column $A_{ij}$ has exactly one +1+1 and one 1-1, so the reduced cost simplifies to the closed-form expression
@@
A BFS is optimal if and only if $\bar{c}_{\ell}\geq 0$ for all nonbasic indices $\ell\in N$, i.e., $r_{ij}\geq 0$ for all $i,j$.
To perform a pivot, we add the index $(i^{*},j^{*})$ to the basis $B$, inducing a unique cycle in the associated flow graph. Then choose the step size $\Delta$ as the largest admissible increase along the fundamental cycle, i.e.,
$\Delta\;=\;\min\{\,x_{e}:e\in\mathcal{C}^{-}\,\},$
where $\mathcal{C}^{-}$ denotes the set of cycle edges on which the flow is decreased. This guarantees that the updated flow remains feasible and nonnegative. Subsequently, update the flow: $x_{i^{*}j^{*}}\leftarrow\Delta$, and for each edge $(u,v)$ along the cycle, the value of $x_{uv}$ increases or decreases by $\Delta$ depending on the direction of the cycle. Finally, the variable that reaches zero and removes it from the basis, completing the pivot step [16]. Each iteration of the NS method involves computing the dual potential vector $\boldsymbol{\pi}$ in $O(M)$ time, evaluating all reduced costs in $O(N)$, and performing pivoting operations in $O(M)$.
@@ 3.2 Block Coordinate Descent Method for Optimal Transport
In this subsection, we introduce the BCD method for the OT problem in the LP form (2).
@@
$D(\mathbf{x};\mathcal{H})\triangleq\arg\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}$. (5)
$q(\mathbf{x};\mathcal{H})\triangleq\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}$. (6)
Here, $D(\mathbf{x};\mathcal{H})$ represents the set of optimal solutions for the subproblem with $\mathcal{H}$, and $q(\mathbf{x};\mathcal{H})$ denotes the optimal objective value corresponding to $d\in D(\mathbf{x};\mathcal{H})$ [15].
@@
The existing BCD method [18], randomly selects block coordinates based on certain properties of specific matrices. In [18], the authors propose the Accelerated Random Block Coordinate Descent algorithm and establish the following convergence theorem.
@@
The gap between the current objective function value and the optimal objective function value converges with rate 1v1-v, i.e.,
@@
These approaches provide a simple mechanism for constructing subproblems by selecting block coordinates. However, when the size n of the OT problems is large, $v$ becomes extremely small, leading to slow convergence. Furthermore, these approaches do not account for the sparsity of the solution of the OT problem. Many variables that ultimately take zero values continue to be included in the block coordinates at each iteration. Moreover, these approaches typically provide only asymptotic convergence guarantees, rather than finite termination at an exact optimal solution.
@@ 3.3 Sinkhorn Algorithm for Optimal Transport
Instead of solving the classical OT problem defined in (1), the Sinkhorn algorithm [6] introduces an entropy regularization term
@@
The regularization parameter $\varepsilon$ introduces a smoothing effect in the OT problem. As $\varepsilon$ decreases, the solution becomes closer to the exact OT plan, but the algorithm may converge more slowly. Conversely, larger values of $\varepsilon$ result in faster computations, but with a higher approximation error in the transport plan.